OpenStudy (ronnie_12):

Find the difference. 4d2e - d2e 3 3de 3d2e 4d2e

rebeccaxhawaii (rebeccaxhawaii):

WELCOME TO OSSSS but you are in the wrong section http://prntscr.com/alrznj then go to correct section but ill help you c:

rebeccaxhawaii (rebeccaxhawaii):

\[4d^2e-d^2e\] right?

OpenStudy (ronnie_12):

Yah

rebeccaxhawaii (rebeccaxhawaii):

well \[d^2e=1d^2e\] so \[4d^2e-1d^2e=?\]

OpenStudy (ronnie_12):

Yes

rebeccaxhawaii (rebeccaxhawaii):

what is it? \[4d^2e-1d^2e=?\]

TheSmartOne (thesmartone):

Hint: \(\bf\Large 4d^2e - d^2e = d^2e(4-1) \) So what is 4 - 1 = ?

rebeccaxhawaii (rebeccaxhawaii):

\[4d^2e-1d^2e=(4-1)d^2e\]

OpenStudy (tennis5518):

Hey there!!! Since you are new here, read this tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 You might not have known, but you are in the "OpenStudy Feedback" section. Place your mouse over where it says OpenStudy Feedback, and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied! Please enjoy being on OpenStudy. And don’t forget to read OpenStudy's Code of Conduct and also, OpenStudy's Terms and Conditions This section is for providing your Feedback/Appreciation regarding the site and by Feedback, we mean "helpful information or criticism that is given to someone to say what can be done to improve a product (which is the site in this case)" You can use this section: ⇝ If you want to appreciate to OpenStudy ⇝ To share your thoughts on how to improve OpenStudy ⇝ To inform the site about any difficulty you are facing while using OpenStudy ⇝ To ask any other question related to this site. ⇝ If you want to post any posts such as congratulating someone, saying Happy Birthday, etc. If you have any questions, please feel free to PM me or tag me in your post And to get you started, let me introduce some great users for you to meet @Wendy.Ivette11714 @YanaSidlinskiy @PillowFace750 @ILovePuppiesLol @rebeccaxhawaii @RhondaSommer @jabez177 @Jaynator495 @poopsiedoodle @pooja195

OpenStudy (pillowface750):

:)

rebeccaxhawaii (rebeccaxhawaii):

c: